Let (\Omega , F, P) be a probability space and let E1 and E2 be two events with P(E₁) = 0.2, P(E₂) 0.4 and P(E₁/E₂)=0.1.


Find the probability that

(a) Exactly one of the events E1 or E₂ will occur.


(b) At least one of the events E1, or E2 will occur.


(c) None of E1 and E2 will occur


1
Expert's answer
2021-08-18T16:35:05-0400

(a)


"P(\text{Exactly one of the events }E_1\ or\ E_2)"

"=P(E_1)-P(E_1\cap E_2)+P(E_2)-P(E_1\cap E_2)"

"=0.2-0.1+0.4-0.1=0.4"

(b)


"P(\text{At least one of the events }E_1\ or\ E_2)"

"=P(E_1\cup E_2)=P(E_1)+P(E_2)-P(E_1\cap E_2)"

"=0.2+0.4-0.1=0.5"

(c)


"P(\text{None of }E_1\ and\ E_2)"

"=1-P(E_1\cup E_2)=1-0.5=0.5"